Soybean meal is 14% protein; cornmeal is 7% protein. How many pounds of each should be mixed together in order to get a 280 pound mixture that is 10% protein
Answer by Earlsdon(6294)

Let x = the number of pounds of the 14% protein Soybean meal, then (280-x) = the number of pounds of the 7% protein Cornmeal.
Write the equation that relates the amount of protein in the two amounts to be mixed to the final mixture of 280 pounds of 10% protein mix.
Change the percentages to their decimal equivalents.
0.14x+0.07(280-x) = 0.1(280) Simplify.
0.14x+19.6-0.07x = 28 Combine the x's.
0.07x+19.6 = 28 Subtract 19.6 from both sides.
0.07x = 8.4 Divide both sides by 0.07
x = 120 and (280-x) = 160
You will need to mix 120 lbs of the 14% protein Soybean meal with 160 lbs of the 7% Cornmeal to obtain 280 lbs of 10% protein mixture.